OpenStudy (anonymous):

Find the percent of the number. 5% of 800 A. 4000 B. 400 C. 40 D. 0.40

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

40 .08*800=40

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Find the percent of the number. 522% of 85 A. 4.437 B. 44.37 C. 443.7 D. 44,370

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

443.7 5.22*85=443.7

OpenStudy (anonymous):

Sorry I was busy.

OpenStudy (anonymous):

72 is 75% of what number? A. 0.96 B. 54 C. 96 D. 5400

OpenStudy (anonymous):

96 .75*96=72

OpenStudy (anonymous):

80% of a number is 96. What is 25% of that same number? A. 20 B. 30 C. 40 D. 120

OpenStudy (anonymous):

I can't figure it out. @horseloverallthetime

OpenStudy (anonymous):

B is the answer :P @Tellme2242